Direct Materials Purchases Budget

Soda is the largest bottler of Soda in Western Europe. The company purchases Brand 1 and Brand 2 concentrate from The Soda Company, dilutes and mixes the concentrate with carbonated water, and then fills the blended beverage into cans or plastic two-liter bottles. Assume that the estimated production for Brand 1 and Brand 2 two-liter bottles at the Wakefield, UK, bottling plant is as follows for the month of May:

Brand 1 146,000 two-liter bottles
Brand 2 111,000 two-liter bottles

In addition, assume that the concentrate costs $72 per pound both Brand 1 and Brand 2 and is used at a rate of 0.15 pound per 100 liters of carbonated water in blending Brand 1 and 0.2 pound per 100 liters of carbonated water in blending Brand 2. Assume that two liters of carbonated water are used for each two-liter bottle of finished product. Assume further that two-liter bottles cost $0.099 per bottle and carbonated water costs $0.079 per liter.

Prepare a direct materials purchases budget for May, assuming inventories are ignored, because there are no changes between beginning and ending inventories for concentrate, bottles, and carbonated water.

Soda-Wakefield Plant
Direct Materials Purchases Budget
For the Month Ending May 31 (assumed data)
Concentrate 2-Liter Bottles Carbonated Water
Materials required for production:
Brand 1 lbs. btls. ltrs.
Brand 2
Total materials lbs. btls. ltrs.
Direct materials unit price x $ x $ x $
Total direct materials to be purchased $ $ $
Answer #1

Materials required for production:

Concentrate:
Brand 1 = (146,000 * 2 ltr) * 0.15 / 100 = 438 lbs
Brand 2 = (111,000 * 2 ltr) * 0.20 / 100 = 444 lbs

Bottles:
Brand 1 = 146,000
Brand 2 = 111,000

Carbonated Water:
Brand 1 = 146,000 * 2 ltr = 292,000 ltrs
Brand 2 = 111,000 * 2 ltr = 222,000 ltrs